OpenStudy (anonymous):

Find the equation of the quadratic function with zeros 16 and 18 and vertex at (17, -2).

OpenStudy (anonymous):

y = a (x - 18) (x - 16) plug (17,-2) in for x and y and solve for a

OpenStudy (anonymous):

do any of yall know this? @mathmale @jim_thompson5910 @ranga

OpenStudy (anonymous):

i got an answer that isnt one of my choices.. im confused @sourwing

OpenStudy (anonymous):

you need to expand the equation if the choices are in standard form. Or you need to put it in vertex form if the choices are vertex form

jimthompson5910 (jim_thompson5910):

as sourwing is saying, y = a (x - 18) (x - 16) y = a (17 - 18) (17 - 16) ... plug in x = 17 -2 = a (17 - 18) (17 - 16) ... plug in y = -2 -2 = a*(-1)*(1) -2 = -a a = 2

jimthompson5910 (jim_thompson5910):

so you now have y = 2(x - 18)(x - 16) expand that out if necessary

OpenStudy (anonymous):

how do i expand?

jimthompson5910 (jim_thompson5910):

create a 2x2 box with the terms outside like so